# Variations of metric that preserve a Riemannian submersion and geometry of its fibers

### v1

@@ Abstract
On the domain of a Riemannian submersion, we consider variations (i.e., smooth one-parameter families) of Riemannian metrics, for which the submersion is Riemannian and which all keep the metric induced on its fibers fixed. We obtain a formula for the variation of the second fundamental form of the fibers with respect to such changes of metric. We find a choice of parameters defining the variations, that allows to easily formulate the necessary and sufficient conditions for preserving particular geometry of the fibers, i.e., keeping them totally geodesic, totally umbilical, or minimal. These conditions are related to the existence of Killing, conformal Killing and divergence-free vector fields on the fibers. We find conditions for metric to be a critical point of integrated squared norms of the mean curvature and the second fundamental form of the fibers, with respect to the considered variations, and prove that all critical points of these functionals are points of their local minima. We also examine variations of sectional curvatures of planes defined by the horizontal lifts of vectors from the image of the submersion.
@@ 1 Introduction
The description of one-parameter families (called variations) of metrics, that preserve a given Riemannian submersion, gives a straightforward condition, relating components of $B_{t}=\partial_{t}g_{t}$ tensor field, given in Theorem 3.2. Obtaining such variation amounts to solving a certain system of linear ordinary differential equations at every point of $M$, where particular components of $B_{t}$ in a $g_{0}$-orthonormal frame are parameters controlling the ODEs and can be set arbitrarily. However, we find it more useful to consider an equivalent, non-linear ODE system determining the metric, with parameters closely related to properties of $B_{t}^{\sharp}$ tensor field, which is defined by $g_{t}(B_{t}^{\sharp}X,Y)=B_{t}(X,Y)$ for all vector fields $X,Y$ on $M$. These new parameters are a set of vertical vector fields, paired with $g_{0}$-horizontal lifts of vector fields from $N$, defined in Corollary 3.6. In Section 3.2, we find a simple formula for variation of the second fundamental form of the fibers, in terms of those vector fields, and conditions for keeping the fibers totally geodesic, totally umbilical or minimal - which require the vector fields to be Killing, conformal Killing or divergence-free, when restricted to the fibers (respectively).
@@
In Section 4 we examine variational problems for extrinsic geometry of the fibers of a Riemannian submersion $\pi$, considering $L_{2}$ norms of the second fundamental form and the mean curvature of the fibers as functionals on the space of those Riemannian metrics, that preserve the Riemannian submersion and are fixed along its fibers. These functionals can be also defined for a single fiber; their critical points, found in Section 4.1, are metrics satisfying simply formulated conditions, respectively: vanishing of a certain divergence-type operator on the second fundamental form, and the mean curvature of the fibers being a projectable vector field. The choice of parameters determining variations allows to easily compute the second variations of the functionals in Section 4.2, which at all critical points are non-negative. In Section 4.3 we examine the sectional curvature of the horizontal lift of a local plane field on $N$ and its variations with respect to the considered changes of metric, that preserve the Riemannian submersion, but move the horizontal lift with respect to the fibers. This variational problem admits as critical points only metrics, where the distribution arising as the horizontal lift of the plane field satisfies a certain integrability condition. In general, variation of such curvature at $x\in M$ depends both on behaviour of vector fields defining the plane on $N$ in a neighbourhood of $\pi(x)$, and on behaviour of vector fields defining the variation in a neighbourhood of $x$. On the other hand, variational problem for integrated horizontal scalar curvature (i.e., the sum of sectional curvatures of planes defined by an orthogonal set of horizontal vectors) on a compact domain of a Riemannian submersion is well posed and reduces to a problem for the norm of the integrability tensor of the horizontal distribution, examined e.g., in [14, 16].
@@ 2 Notation and definitions
The distribution ${\cal V}=\ker\pi_{*}$ is called vertical, its $g$-ortogonal complement ${\cal H}$ is called the horizontal distribution; a vector field is called vertical or horizontal if it has values in the corresponding distribution. We will use notation $n=\dim\ker\pi_{*}$ and $n+p=\dim M$. We say that a vector field $X\in\mathfrak{X}_{M}$ is projectable if $\pi_{*}X$ is well defined; $X$ is projectable if and only if $[X,V]\in\mathfrak{X}_{\ker\pi_{*}}$ for all $V\in\mathfrak{X}_{\ker\pi_{*}}$. For every $W\in\mathfrak{X}_{N}$ there exists a $g$-horizontal lift of $W$, i.e., a unique $g$-horizontal vector field $\pi^{*}W$ on $M$ such that $\pi_{*}(\pi^{*}W)=W$. For all $x\in N$, let ${\cal F}_{x}=\pi^{-1}(\{x\})$ be the fiber of $\pi$ over $x$.
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion. For a one-parameter family $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$, let ${\cal H}(t)$ denote the horizontal distribution of $\pi$ on $(M,g_{t})$, i.e., the $g_{t}$-orthogonal complement of $\ker\pi_{*}$. Let $P_{{\cal V}}^{t}$ and $P_{{\cal H}}^{t}$ denote the $g_{t}$-orthogonal projections onto the distributions $\ker\pi_{*}$ and ${\cal H}(t)$, respectively. We shall consider sets of linearly independent vector fields $\{E_{1},\ldots,E_{n},{\cal E}_{n+1},\ldots{\cal E}_{n+p}\}$ defined on an open subset of $M$, where $E_{a}\in\mathfrak{X}_{\ker\pi_{*}}$ for all $a\in\{1,\ldots,n\}$; we will call them local adapted frames and denote such a frame shortly by $\{E_{a},{\cal E}_{i}\}$. If vector fields in a frame $\{E_{a},{\cal E}_{i}\}$ are $g$-orthonormal, we will call such frame $g$-orthonormal. For indices we use the convention $a,b,c\in\{1,\ldots,n\}$, $i,j,k\in\{n+1,\ldots,n+p\}$ and $\mu,\nu\in\{1,\ldots,n+p\}$.
@@
for all $X,Y\in\mathfrak{X}_{{\cal H}(t)}$ and all $Z\in\ker\pi_{*}$. If $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion, we have ${\tilde{h}}_{t}=0$ and ${\tilde{A}}^{t}=0$ [13]. The second fundamental form of the fibers of $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is defined as
@@
for all $X,Y\in\mathfrak{X}_{\ker\pi_{*}}$; we have $h_{t}(X,Y)=h_{t}(Y,X)$. We say that the fibers of $\pi:(M,g_{t})\rightarrow(N,g_{N})$ are totally geodesic if for all $X,Y\in\ker\pi_{*}$ we have $h_{t}(X,Y)=0$. Let $\{E_{1},\ldots,E_{n}\}$ be a local $g_{t}$-orthonormal frame spanning $\ker\pi_{*}$, we define the mean curvature of the fibers as
@@
We say that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ has minimal fibers if $H_{t}=0$. We say that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ has totally umbilical fibers if for all $X,Y\in\mathfrak{X}_{\ker\pi_{*}}$ we have $h_{t}(X,Y)=\frac{1}{n}g_{t}(X,Y)H_{t}$.
@@
$\operatorname{div}^{t}_{\ker\pi_{*}}X=\sum\nolimits_{a=1}^{n}g_{t}(\nabla^{t}_{E_{a}}X,E_{a}),$
where $\{E_{1},\ldots,E_{n}\}$ are local $g_{t}$-orthonormal vertical vector fields. Let $X\in\mathfrak{X}_{{\cal F}_{x}}$, then $\operatorname{div}^{t}_{\ker\pi_{*}}X$ defined by the above formula is the divergence of $X$ on $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ and if $\operatorname{div}^{t}_{\ker\pi_{*}}X=0$, we will say that such $X$ is divergence-free.
@@
for all $X\in\mathfrak{X}_{{\cal F}_{x}}$ and $S\in{\cal S}_{r,2}({\cal F}_{x})$. We note that for $r=0$, ${\cal T}_{r,1}({\cal F}_{x})=\Omega^{1}({\cal F}_{x})$ is the space of $1$-forms on ${\cal F}_{x}$ and the formal adjoint of $\delta_{t}$ is given by $(\delta^{*}_{t}\omega)(X,Y)=\frac{1}{2}(\mathcal{L}_{\omega^{\sharp}}g_{t})(X,Y)$, for all $\omega\in\Omega^{1}({\cal F}_{x})$ and all $X,Y\in\mathfrak{X}_{{\cal F}_{x}}$, where $\mathcal{L}_{\omega^{\sharp}}g_{t}$ is the Lie derivative of $g_{t}$ with respect to vector field $\omega^{\sharp}$, which is defined by $g_{t}(\omega^{\sharp},Z)=\omega(Z)$ for all $Z\in\mathfrak{X}_{{\cal F}_{x}}$ [3].
@@ 3.1 Preserving Riemannian submersions
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion. We say that a variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ preserves the Riemannian submersion $\pi$, if $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$.
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion. A variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ preserves the Riemannian submersion $\pi$ if and only if for all $t\in(-\epsilon,\epsilon)$
@@
Using $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$, we obtain
$B_{t}(U,V)=0\quad{\rm for\;all}\;U,V\in\mathfrak{X}_{\ker\pi_{*}}$ (2)
@@
The above holds if and only if for all $x\in M$, $X,Y\in T_{x}M$ and every $g_{0}$-orthonormal – and hence $g_{t}$-orthonormal for all $t\in(-\epsilon,\epsilon)$, because $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ – basis $\{E_{1},\ldots,E_{n}\}$ of ${\ker\pi_{*}}_{x}$ we have
00 0 $=$ $\partial_{t}g_{t}(P_{{\cal H}}^{t}X,P_{{\cal H}}^{t}Y)=\partial_{t}\big{(}g_{t}(X-\sum\nolimits_{a=1}^{n}g_{t}(X,E_{a})E_{a}\,,Y-\sum\nolimits_{a=1}^{n}g_{t}(Y,E_{b})E_{b})\big{)}$ (3)
@@
$+\partial_{t}\sum\nolimits_{a=1}^{n}g_{t}(X,E_{a})g_{t}(E_{a},Y)$
$=$ $B_{t}(X,Y)-\sum\nolimits_{a=1}^{n}B_{t}(X,E_{a})g_{t}(Y,E_{a})-\sum\nolimits_{a=1}^{n}B_{t}(Y,E_{a})g_{t}(X,E_{a}).$
From (3), $P_{{\cal V}}^{t}Z=\sum\nolimits_{a=1}^{n}g_{t}(Z,E_{a})E_{a}$, $Z=P_{{\cal V}}^{t}Z+P_{{\cal H}}^{t}Z$ for all $Z\in\mathfrak{X}_{M}$ and (2), we obtain (1). ∎
In particular, from (1) follows that every variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$, that preserves the Riemannian submersion $\pi$, is determined by values of $B_{t}(Z,V)$ for all $Z\in\mathfrak{X}_{M}$ and $V\in\mathfrak{X}_{\ker\pi_{*}}$.
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion. If a variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ preserves the Riemannian submersion $\pi$, then in every local adapted $g_{0}$-orthonormal frame $\{E_{a},{\cal E}_{i}\}$ components $g_{ab}(t)=g_{t}(E_{a},E_{b})$, $g_{ai}(t)=g_{t}(E_{a},{\cal E}_{i})$, $g_{ij}(t)=g_{t}({\cal E}_{i},{\cal E}_{j})$ of $g_{t}$ are the solution of the following system of ODEs:
@@
$\partial_{t}g_{ij}=\sum\nolimits_{a=1}^{n}\left(b_{ia}g_{aj}+b_{ja}g_{ai}\right),$
@@
On the other hand, for an adapted $g_{0}$-orthonormal frame $\{E_{a},{\cal E}_{i}\}$ on an open, relatively compact set $U\subset M$, every solution of (4)-(6), with initial conditions $g_{ai}(0)=0$ and $g_{ij}(0)=\delta_{ij}$ and arbitrary smooth functions $\{b_{ai}\}$, determines some variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(U,\ker\pi_{*},g_{0})$, that preserves the Riemannian submersion $\pi$.
@@
For an adapted $g_{0}$-orthonormal frame $\{E_{a},{\cal E}_{i}\}$, equation (4) is equivalent to (2) and $\{E_{a}\}$ being $g_{0}$-orthonormal. Equation (6) is equivalent to $\partial_{t}g_{t}(E_{a},{\cal E}_{i})=B_{t}(E_{a},{\cal E}_{i})$, and (6) is equivalent to (3), written for $X={\cal E}_{i}$, $Y={\cal E}_{j}$.
On the other hand, for every $x\in U$ there exists a neighbourhood on which for arbitrary smooth functions $\{b_{ai}\}$, the solution $\{g_{\mu\nu}(t)\}$ of (4)-(6) with $g_{ai}(0)=0$ and $g_{ij}(0)=\delta_{ij}$ exists on interval $t\in(-\epsilon,\epsilon)$ for some $\epsilon>0$. By decreasing $\epsilon$, if necessary, we can assume that $\{g_{\mu\nu}(t)\}$ are components of positive-definite matrix for all $t\in(-\epsilon,\epsilon)$, since $g_{\mu\nu}(0)=\delta_{\mu\nu}$. From the compactness of ${\overline{U}}$ it follows, that there exists $\epsilon>0$ such that solution of these properties exists for all $t\in(-\epsilon,\epsilon)$ on $U$. Hence, $\{g_{\mu\nu}(t)\}$ define a Riemannian metric on $U$, determined by its components in the frame $\{E_{a},{\cal E}_{i}\}$ and bilinearity. ∎
Theorem 3.4.
Let $\{E_{a},{\cal E}_{i}\}$ be a local adapted $g_{0}$-orthonormal frame. Then $\{g_{\mu\nu}(t)\}$ is the solution of the system (4)-(6), with $g_{\mu\nu}(0)=\delta_{\mu\nu}$, on some interval $t\in(-\epsilon,\epsilon)$ if and only if $\{g_{\mu\nu}(t)\}$ is the solution of the following system:
@@
with $g_{\mu\nu}(0)=\delta_{\mu\nu}$, on some interval $t\in(-\epsilon^{\prime},\epsilon^{\prime})$.
@@
Indeed, (9)-(9) is a particular case of (4)-(6), with
@@
Thus, every solution of (9)-(9) is a solution of (4)-(6) with $\{b_{ai}\}$ as in (10).
On the other hand, let $g_{t}$ be the Riemannian metric determined by a solution of (4)-(6) for $t\in(-\epsilon,\epsilon)$. From (4) we obtain $g_{t}(B_{t}^{\sharp}E_{a},E_{b})=0$ for all $a,b\in\{1,\ldots,n\}$ and $t\in(-\epsilon,\epsilon)$, hence, $B_{t}^{\sharp}E_{a}\in{\cal H}(t)$. There exists $\epsilon^{\prime}>0$ such that for all $t\in(-\epsilon^{\prime},\epsilon^{\prime})$ vector fields $\{P_{{\cal H}}^{t}{\cal E}_{i}\}$ span ${\cal H}(t)$ and therefore
$B^{\sharp}_{t}E_{a}=\sum\nolimits_{i=n+1}^{n+p}\left(\lambda_{ai}{\cal E}_{i}-\sum\nolimits_{c=1}^{n}g_{ic}\lambda_{ai}E_{c}\right)$ (11)
for some functions $\lambda_{ai}$. For $B_{t}^{\sharp}E_{a}$ given by (11), from $b_{ai}=B_{t}(E_{a},{\cal E}_{i})=g_{t}(B^{\sharp}_{t}E_{a},{\cal E}_{i})$ we obtain that (10) holds, and hence $g_{t}$ is the solution of (9)-(9). ∎

### v2

@@ Abstract
On the domain of a Riemannian submersion, we consider variations (i.e., smooth one-parameter families) of Riemannian metrics, for which the submersion is Riemannian and which all keep the metric induced on its fibers fixed. We obtain a formula for the variation of the second fundamental form of the fibers with respect to such changes of metric. We find a choice of parameters defining the variations, that allows to easily formulate the necessary and sufficient conditions for preserving particular geometry of the fibers, i.e., keeping them totally geodesic, totally umbilical, or minimal. These conditions are related to the existence of Killing, conformal Killing and divergence-free vector fields on the fibers. We find conditions for metric to be a critical point of integrated squared norms of the mean curvature and the second fundamental form of the fibers, with respect to the considered variations, and prove that at all critical points of these functionals the second variation is non-negative. We also examine variations of sectional curvatures of planes defined by the horizontal lifts of vectors from the image of the submersion. In particular, we find that some variations preserving Riemannian submersions with totally geodesic fibers can make vertizontal curvatures not constant on the fibers.
@@ 1 Introduction
The description of one-parameter families (called variations) of metrics, that preserve a given Riemannian submersion, gives a straightforward condition, relating components of $B_{t}=\partial_{t}g_{t}$ tensor field, given in Theorem 3.2. Obtaining such variation amounts to solving a certain system of linear ordinary differential equations at every point of $M$, where particular components of $B_{t}$ in a $g_{0}$-orthonormal frame are parameters controlling the ODEs and can be set arbitrarily. However, we find it more useful to consider an equivalent, non-linear ODE system determining the metric, with parameters closely related to properties of $B_{t}^{\sharp}$ tensor field, which is defined by $g_{t}(B_{t}^{\sharp}X,Y)=B_{t}(X,Y)$ for all vector fields $X,Y$ on $M$. These new parameters are a set of vertical vector fields, paired with $g_{0}$-horizontal lifts of vector fields from $N$, defined in Theorem 3.6. In Section 3.2, we find a simple formula for variation of the second fundamental form of the fibers, in terms of those vector fields, and conditions for keeping the fibers totally geodesic, totally umbilical or minimal - which require the vector fields to be Killing, conformal Killing or divergence-free (respectively), when restricted to the fibers.
@@
In Section 4 we examine variational problems for extrinsic geometry of the fibers of a Riemannian submersion $\pi$, considering $L_{2}$ norms of the second fundamental form and the mean curvature of the fibers as functionals on the space of those Riemannian metrics, that preserve the Riemannian submersion and are fixed along its fibers. These functionals can be also defined for a single fiber; their critical points, found in Section 4.1, are metrics satisfying simply formulated conditions, respectively: vanishing of a certain divergence-type operator on the second fundamental form, and the mean curvature of the fibers being a projectable vector field. The choice of parameters determining variations allows to easily compute the second variations of the functionals in Section 4.2, which at all critical points are non-negative.
In Section 5 we consider how the sectional curvatures in horizontal and vertical directions change with variations of metric preserving a Riemannian submersion. In Section 5.1 we examine the sectional curvature of the horizontal lift of a local plane field on $N$ and its variations with respect to the considered changes of metric, that preserve the Riemannian submersion, but move the horizontal lift with respect to the fibers. This variational problem admits as critical points only metrics for which the distribution arising as the horizontal lift of the plane field satisfies a certain integrability condition. In general, variation of such curvature at $x\in M$ depends both on behaviour of vector fields defining the plane on $N$ in a neighbourhood of $\pi(x)$, and on behaviour of vector fields defining the variation in a neighbourhood of $x$. On the other hand, variational problem for integrated horizontal scalar curvature (i.e., the sum of sectional curvatures of planes defined by an orthogonal set of horizontal vectors) on a compact domain of a Riemannian submersion is well posed and reduces to a problem for the norm of the integrability tensor of the horizontal distribution, examined e.g., in [16, 18].
In Section 5.2 we consider variations of the vertizontal curvatures, i.e., sectional curvatures of planes spanned by a horizontal and a vertical vector. Assuming existence of certain Killing fields on the fibers, in Corollary 5.7 we demonstrate existence of Riemannian submersions with totally geodesic fibers and vertizontal curvatures non-constant on a fiber. This construction can be applied to modify e.g., the Hopf fibrations by $3$-spheres, producing new fat Riemannian submersions [20].
@@ 2 Notation and definitions
[…]
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion. For a one-parameter family $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$, let ${\cal H}(t)$ denote the $g_{t}$-horizontal distribution of $\pi$. Let $P_{{\cal V}}^{t}$ and $P_{{\cal H}}^{t}$ denote the $g_{t}$-orthogonal projections onto the distributions $\cal V$} and ${\cal H}(t)$, respectively. We shall consider sets of linearly independent vector fields $\{E_{1},\ldots,E_{n},{\cal E}_{n+1},\ldots{\cal E}_{n+p}\}$ defined on an open subset of $M$, where $E_{a}\in\mathfrak{X}_{{\cal V}}$ for all $a\in\{1,\ldots,n\}$; we will call them local adapted frames and denote such a frame shortly by $\{E_{a},{\cal E}_{i}\}$. If vector fields in a frame $\{E_{a},{\cal E}_{i}\}$ are $g$-orthonormal, we will call such frame $g$-orthonormal. For indices we use the convention $a,b,c\in\{1,\ldots,n\}$, $i,j,k\in\{n+1,\ldots,n+p\}$ and $\mu,\nu\in\{1,\ldots,n+p\}$.
@@
for all $X,Y\in\mathfrak{X}_{{\cal H}(t)}$ and all $Z\in{\cal V}$. If $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion, we have ${\tilde{h}}_{t}=0$ and ${\tilde{A}}^{t}=0$ [15]. The second fundamental form of the fibers of $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is defined as
@@
for all $X,Y\in\mathfrak{X}_{{\cal V}}$; we have $h_{t}(X,Y)=h_{t}(Y,X)$. We say that the fibers of $\pi:(M,g_{t})\rightarrow(N,g_{N})$ are totally geodesic if for all $X,Y\in\mathfrak{X}_{{\cal V}}$ we have $h_{t}(X,Y)=0$. Let $\{E_{1},\ldots,E_{n}\}$ be a local $g_{t}$-orthonormal frame spanning $\cal V$}, we define the mean curvature of the fibers as
@@
We say that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ has minimal fibers if $H_{t}=0$. We say that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ has totally umbilical fibers if for all $X,Y\in\mathfrak{X}_{{\cal V}}$ we have $h_{t}(X,Y)=\frac{1}{n}g_{t}(X,Y)H_{t}$.
@@
$\operatorname{div}^{t}_{{\cal V}}X=\sum\nolimits_{a=1}^{n}g_{t}(\nabla^{t}_{E_{a}}X,E_{a}),$
where $\{E_{1},\ldots,E_{n}\}$ are local $g_{t}$-orthonormal vertical vector fields. Let $X\in\mathfrak{X}_{{\cal F}_{x}}$, then $\operatorname{div}^{t}_{{\cal V}}X$ defined by the above formula is the divergence of $X$ on the Riemannian manifold $({\cal F}_{x},g_{t}|_{{\cal F}_{x}})$ and if $\operatorname{div}^{t}_{{\cal V}}X=0$, we will say that such $X$ is divergence-free on $({\cal F}_{x},g_{t}|_{{\cal F}_{x}})$.
@@
for all $X\in\mathfrak{X}_{{\cal F}_{x}}$ and $S\in{\cal S}_{r,2}({\cal F}_{x})$. We note that for $r=0$, ${\cal T}_{0,1}({\cal F}_{x})=\Omega^{1}({\cal F}_{x})$ is the space of $1$-forms on ${\cal F}_{x}$ and the formal adjoint of $\delta_{t}$ is given by $(\delta^{*}_{t}\omega)(X,Y)=\frac{1}{2}(\mathcal{L}_{\omega^{\sharp}}g_{t})(X,Y)$, for all $\omega\in\Omega^{1}({\cal F}_{x})$ and all $X,Y\in\mathfrak{X}_{{\cal F}_{x}}$, where $\mathcal{L}_{\omega^{\sharp}}g_{t}$ is the Lie derivative of $g_{t}$ with respect to vector field $\omega^{\sharp}$, which is defined by $g_{t}(\omega^{\sharp},Z)=\omega(Z)$ for all $Z\in\mathfrak{X}_{{\cal F}_{x}}$ [3].
@@ 3.1 Preserving Riemannian submersions
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion. We say that a variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ preserves the Riemannian submersion $\pi$, if $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$.
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion. A variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ preserves the Riemannian submersion $\pi$ if and only if for all $t\in(-\epsilon,\epsilon)$
@@
Using $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$, we obtain
$B_{t}(U,V)=0\quad{\rm for\;all}\;U,V\in\mathfrak{X}_{{\cal V}}$ (2)
@@
The above holds if and only if for all $x\in M$, $X,Y\in T_{x}M$ and every $g_{0}$-orthonormal – and hence $g_{t}$-orthonormal for all $t\in(-\epsilon,\epsilon)$, because $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ – basis $\{E_{1},\ldots,E_{n}\}$ of ${{\cal V}}_{x}$ we have
$0$ $=$ $\partial_{t}g_{t}(P_{{\cal H}}^{t}X,P_{{\cal H}}^{t}Y)=\partial_{t}\big{(}g_{t}(X-\sum\nolimits_{a=1}^{n}g_{t}(X,E_{a})E_{a}\,,Y-\sum\nolimits_{a=1}^{n}g_{t}(Y,E_{b})E_{b})\big{)}$ (3)
@@
$+\partial_{t}\sum\nolimits_{a,b=1}^{n}g_{t}(X,E_{a})g_{t}(E_{b},Y)g_{t}(E_{a},E_{b})$
$=$ $B_{t}(X,Y)-\sum\nolimits_{a=1}^{n}B_{t}(X,E_{a})g_{t}(Y,E_{a})-\sum\nolimits_{a=1}^{n}B_{t}(Y,E_{a})g_{t}(X,E_{a}),$
where we used $g_{t}(E_{a},E_{b})=g_{0}(E_{a},E_{b})=\delta_{ab}$. From (3), $P_{{\cal V}}^{t}Z=\sum\nolimits_{a=1}^{n}g_{t}(Z,E_{a})E_{a}$, $Z=P_{{\cal V}}^{t}Z+P_{{\cal H}}^{t}Z$ for all $Z\in\mathfrak{X}_{M}$ and (2), we obtain (1). ∎
In particular, from (1) follows that every variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$, that preserves the Riemannian submersion $\pi$, is determined by values of $B_{t}(Z,V)$ for all $Z\in\mathfrak{X}_{M}$ and $V\in\mathfrak{X}_{{\cal V}}$.
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion. A variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ preserves the Riemannian submersion $\pi$ if and only if in every local adapted $g_{0}$-orthonormal frame $\{E_{a},{\cal E}_{i}\}$ components $g_{ab}(t)=g_{t}(E_{a},E_{b})$, $g_{ai}(t)=g_{t}(E_{a},{\cal E}_{i})$, $g_{ij}(t)=g_{t}({\cal E}_{i},{\cal E}_{j})$ of $g_{t}$ are the solution, on interval $t\in(-\epsilon,\epsilon)$, of the following system:
@@
$\partial_{t}g_{ij}=\sum\nolimits_{a=1}^{n}\left(b_{ia}g_{aj}+b_{ja}g_{ai}\right),$ (6)
@@
For an adapted $g_{0}$-orthonormal frame $\{E_{a},{\cal E}_{i}\}$, equation (4) is equivalent to (2) and $\{E_{a}\}$ being $g_{0}$-orthonormal. Equation (5) is equivalent to $\partial_{t}g_{t}(E_{a},{\cal E}_{i})=B_{t}(E_{a},{\cal E}_{i})$, and (6) is equivalent to (3), written for $X={\cal E}_{i}$, $Y={\cal E}_{j}$. We note that, due to (2), equation (3) written for $X={\cal E}_{i}$ and $Y=E_{b}$ yields (5), i.e., it does not impose any conditions on $\{b_{ai}\}$. ∎
Lemma 3.4.
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion. Let $W\subset M$ be an open set on which there exists an adapted $g_{0}$-orthonormal frame $\{E_{a},{\cal E}_{i}\}$. Let $\{\lambda_{ai}:W\rightarrow\mathbb{R}\}$ be smooth functions. Let $U$ be an open, relatively compact set, such that $\overline{U}\subset W$. Then there exists a variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(U,{\cal V},g_{0})$, that preserves the Riemannian submersion $\pi$, such that $\{g_{\mu\nu}(t)\}$ is the solution of the following system
@@
with $g_{\mu\nu}(0)=\delta_{\mu\nu}$, and $\{P_{{\cal H}}^{t}{\cal E}_{n+1},\ldots,P_{{\cal H}}^{t}{\cal E}_{n+p}\}$ are linearly independent for all $t\in(-\epsilon,\epsilon)$.
@@
We consider (9)-(9) as a system of ODEs, smoothly depending on parameter $x\in W$. It follows that for every $x\in W$ there exists a neighbourhood ${\cal N}_{x}$ on which the solution $\{g_{\mu\nu}(t)\}$ of (9)-(9) with $g_{ai}(0)=0$ and $g_{ij}(0)=\delta_{ij}$ exists on interval $t\in(-\epsilon_{x},\epsilon_{x})$ for some $\epsilon_{x}>0$. By decreasing $\epsilon_{x}$, if necessary, we can assume that $\{g_{\mu\nu}(t)\}$ are components of positive-definite matrix for all $t\in(-\epsilon_{x},\epsilon_{x})$, because $g_{\mu\nu}(0)=\delta_{\mu\nu}$, and, since $\{{\cal E}_{n+1},\ldots,{\cal E}_{n+p}\}$ are linearly independent, $\{{\cal E}_{i}-\sum\nolimits_{a=1}^{n}g_{ia}(t)E_{a},i=n+1,\ldots,n+p\}$ are linearly independent on ${\cal N}_{x}$ for all $t\in(-\epsilon_{x},\epsilon_{x})$. Then $\{{\cal N}_{x},x\in W\}$ is an open cover of the compact set $\overline{U}$}, so it has a finite subcover and hence there exists $\epsilon>0$ such that solution $\{g_{\mu\nu}(t)\}$ with all above properties exists for all $t\in(-\epsilon,\epsilon)$ on $U$. Hence, $\{g_{\mu\nu}(t)\}$ define a Riemannian metric $g_{t}$ on $U$, determined by its components in the frame $\{E_{a},{\cal E}_{i}\}$ and bilinearity.
To prove that $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(U,{\cal V},g_{0})$ preserves the Riemannian submersion $\pi$, observe that (2) holds by (9) and (9)-(9) is a particular case of (4)-(6), with
@@
Hence, (3) holds for all vector fields from the adapted frame $\{E_{a},{\cal E}_{i}\}$, and hence for all vector fields on $U$. The last claim follows from $P_{{\cal H}}^{t}{\cal E}_{i}={\cal E}_{i}-\sum\nolimits_{a=1}^{n}g_{ia}(t)E_{a}$. ∎